OpenStudy (accidentalaichan):

Which ratio forms a proportion with 15/ 21?

OpenStudy (accidentalaichan):

\[\frac{ 20 }{ 26 }\] \[\frac{ 7 }{ 5 }\] \[\frac{ 20 }{ 28 }\] \[\frac{ 12 }{ 18 }\]

OpenStudy (anonymous):

15/3=5 21/3=7 So is?

OpenStudy (accidentalaichan):

7/5?

OpenStudy (anonymous):

I don't know why is it 7/5 it's 5/7

OpenStudy (accidentalaichan):

That's why I'm confused

OpenStudy (accidentalaichan):

Unless it's 20/28 because 20 is a multiple of 5 and 28 is a multiple of 7

OpenStudy (anonymous):

20/4=5 28/4=7 :)

OpenStudy (anonymous):

yeah it's 20/28

OpenStudy (accidentalaichan):

Okay, makes sense now. Thank you so much!

OpenStudy (anonymous):

welcome :)